Hello I am trying to figure out how to write this equation in my sons homework I worked the answer in my head 150 but need help showing the work thanks.
David and Amrita had an equal number of marbles. After Amrita gave 50 marbles to David, he had 5 times as many marbles as her. Find the total number of marbles they had.

Answer by Alan3354(69443)

David and Amrita had an equal number of marbles. After Amrita gave 50 marbles to David, he had 5 times as many marbles as her. Find the total number of marbles they had.
--------------------
b = # of marbles they each had
b+50 = 5*(b-50)
b+50 = 5b - 250
-6b = -300
b = 50